OpenStudy (anonymous):

The sum of the squares of two consecutive negative integer is 61. Find the smaller of the two integers. Help pls

OpenStudy (anonymous):

let x equal the smaller of the 2 integers then the next consecutive negative integer would be "x+1" we know the square of the 2 numbers is equal to 61 (x)^2+(x+1)^2=61 just solve for x (remember that x needs to be negative)

OpenStudy (radar):

Are you able to proceed?

OpenStudy (anonymous):

Well I got 2x^2+4x-57=0 After I have foiled and added now I'm stuck :/

OpenStudy (anonymous):

What do I do?

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

You need to redo your foil tho

OpenStudy (radar):

Review your result of (x+1)^2

OpenStudy (kropot72):

The quadratic that you need to solve is this: \[2n ^{2}-2n-60=0\]

OpenStudy (anonymous):

Lol yeah guys sorry i just checked my work your right :)

OpenStudy (radar):

Yes, that is what you are striving for. You can simplify by dividing all terms by 2 getting\[n ^{2}+n - 30 = 0\]after correcting the sign of 2n in kropot72 post.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You can use the quadratic, but if you recognize it, this factors quite nicely

OpenStudy (radar):

Now factor and solve.

OpenStudy (radar):

Have you factored it yet?

OpenStudy (radar):

Yes, I see that the question is now closed. Completeidiot gave excellent guidance. We all tried, I hope some understanding was imparted.